In the Department of Natural Sciences, 14 faculty members have a PhD, and 30 faculty
members do not have a PhD. In the Department, the number of female faculty who do not
have a PhD is 10 more than the number of females who have a PhD. If a third of the male
faculty in the Department have a PhD, then what is the number of female faculty in the
Department with a PhD?
1
Expert's answer
2020-06-14T18:17:34-0400


Let X1 be the number of women with a degree.


X10 -, without degree


Y1 number of men with a degree


Y10 - no degree


Then X1 + Y1 = 14


X10 + Y10 = 30


X10 -X1 + 10


Y10 = 2 * Y1


we have 4 equations with four unknowns.


substitute the last two equations into the second equation


X1 + 2 + Y1 = 20


we subtract the first from here, we get Y1 = 6


therefore X1 = 8